# Emergent Dexterity via Diverse Resets and Large-Scale Reinforcement Learning

### 16 Mar 2026 v1

@@ 4 Simulation Experiments
We aim to address the following questions experimentally - (Q1) Does OmniResetoutperform baselines and scale to tasks beyond current methods? (Q2) How do our key design decisions affect performance? (Q3) Can our experts be used to generate diverse data for sim-to-real transfer?
@@ 4.1 Task Descriptions
4.2 Simulation Expert Baseline Comparisons
We compare to the following baselines, which bootstrap learning with expert demonstrations, providing them more prior information about how to solve the task than OmniReset. For each of the methods, we supply 100 demonstration generated by the final OmniReset policy, effectively giving the baselines access to the optimal behavior. The initial conditions for these demonstrations are drawn from the reaching region $S^{R}$, which corresponds to the set of initial conditions the robot will encounter when solving the full task.
@@ 4.2 Simulation Expert Baseline Comparisons
  • 2) DeepMimic: We use DeepMimic-style reward augmentation (Peng et al., 2018) on top of our generic rewards. During resets, a random demonstration is chosen, and the agent is reset from a random point along the demonstration and is given an auxiliary reward which provides bonuses for following the demonstration.
@@
  • 3) Demo Curriculum: This baseline is constructed in the spirit of method such as (Bauza et al., 2025) and uses a success-weighted autocurriculum to samlple reset states from the demonstrations. We use PPO as the base RL algorithm to ensure a fair comparison.
@@
Learning Curves and Success Rates: We plot learning curves for the Hard variants of each task and the Easy variants of the Peg Insertion, Leg Twisting, and Drawer Insertion tasks in Figure 8. Success rates are reported for initial conditions sampled from the Reaching resets $S^{R}$, which correspond to states from which the robot must solve the full task. We see that OmniReset is able to consistently obtain high success rates on each of the tasks, substantially outperforming baselines. While the baselines are able to make some progress on the Easy variants of the tasks, they consistently struggle to scale to the wider distributions of initial conditions which defines the Hard variants. Figure 5 provides additional insight into the failure modes of baselines. Here we plot success rates during training for initial conditions drawn from the demonstrations which fall into the Near-Goal and Reaching regions of the state-space. This plot demonstrates how the baselines are able to solve part of the task (i.e. when starting near the goal) but fail to scale to the full long-horizon task (i.e. starting from a reaching position).
@@
Robustness of Policies: We conduct a robustness analysis on the learned policies (Fig 10). We sample an initial condition from one of the demonstrations, perturb the initial condition with forces of different magnitudes and report policy success from these perturbed initial conditions. We find that baseline performance quickly degrades under small perturbations, while the performance of OmniReset is barely affected under large perturbations.
@@
We also analyze the ability of policies to solve the task from a wide range of initial conditions in the scatter plots in Figure (Fig 9). These plots show success rates over various initial conditions for both OmniResetand Demo Curriculum (the most successful baseline) on the Leg Screw Easy task. For each plot we show the success rate over 1000 sampled initial conditions from the full task distribution. This plot demonstrates how the baseline struggles to achieve achieve consistent success across the distribution of initial conditions it was trained on, while OmniResetachieves is able to succeed across the entire workspace.
@@ 4.3 Ablating Key Design Decisions
Figure 6: Ablation on grasp sampling range. For this ablation on the screwing task, we find that training RL on narrower grasp sampling ranges leads to worse sample efficiency and lower converged success rate.
@@
Figure 7: Ablation on number of environments. We plot the success rates over course of RL training using different number of environments. We find that the number of environments significantly impacts training performance.
Figure 9: Success states of RL policy. For the screwing task, we plot the xy configurations from which RL polices trained with Demo Curriculum and OmniResetsucceed when trained on the full reset distribution and on a narrow reset distribution. We find that OmniResetsucceeds from a much broader range of initial conditions.
@@ 5 Distillation and Real-World Transfer
We demonstrate the utility of our learned data-generation policies by distilling them into visuomotor policies deployable directly on hardware from RGB inputs. The robot observes $224\times 224$ images from three cameras (front, side, wrist). Using the photorealistic rendering capabilities of IsaacLab (Mittal et al., 2023), we collect 10,000 expert rollouts with synchronized images and actions for standard student–teacher distillation Chen et al. (2021). The student policy uses an ImageNet-pretrained ResNet-18 encoder and a Gaussian MLP head conditioned on the five most recent observations.
@@
Visual randomization. To mitigate the sim-to-real visual gap, we apply extensive domain randomization following DextrAH-G (Singh et al., 2025), varying lighting, backgrounds, object and robot appearance, and workspace textures. Camera extrinsics are calibrated to the real setup with additional pose and FOV jitter for robustness. We also apply standard image augmentations including color jitter, blur, grayscale, and noise.
@@
Dynamics randomization. To reduce the control gap, we calibrate kinematics to hardware and deploy the same task-space operational space controller in simulation and reality, with the policy predicting end-effector pose deltas that are converted to torques using identical Jacobian-based control. We perform system identification of key actuator parameters (e.g., friction and delay) to match real joint behavior, then randomize controller gains and physical parameters around identified values during RL training with a stability curriculum. We additionally apply an action-scaling curriculum to encourage smaller motions and randomize object mass and friction to promote contact robustness.
Real-World Transfer. We successfully deploy a distilled OmniResetpolicy for the Peg Insertion Task, as depicted in Figure 1. Evaluating our distilled RGB policy trained on 10,000 simulation trajectories zero-shot in real achieve a success rate of $25\%$ from a wide range of initial conditions. We compare this to a Diffusion Policy Chi et al. (2025) trained on 100 demonstrations, which achieves only a $4\%$ success rate. Qualitatively, the OmniReset visuomotor policy displays robust retrying behavior and is able to recover and successfully insert the peg even after initial failed attempts caused by the sim-to-real gap. Altogether, this demonstrates how OmniReset provides a foundation for scalable data generation for robust sim-to-real policies which are able to solve tasks from much wider ranges of initial conditions than existing methods.
@@ 6 Conclusion
In this work we presented OmniReset, a simple and scalable system for data generation in simulation for complex, dexterous tasks. The primary insight in OmniResetis showing that a diverse, minimally structured set of reset states paired with large-batch on-policy reinforcement learning in simulation can lead to the emergence of surprisingly complex dexterous behavior. We provide a general purpose recipe to instantiate data generators across a variety of manipulation tasks, and demonstrate both the efficacy of this paradigm in simulation and it’s ability to bootstrap real world policy learning. The presented work is currently still single-task and object, making a multi-task extension in future work a promising direction of investigation. Furthermore, we are excited to extend OmniReset to higher-dimensional settings such as dexterous hand using existing techniques for generating grasps Xu et al. (2023). Furthermore, OmniResetopens up the possibility of more systematically studying scaling laws for policy learning.

### 02 Apr 2026 v3

@@ 4 Simulation Experiments
Let us first consider a study of data generation in simulation with OmniReset. We aim to address the following questions experimentally - (Q1) Does OmniReset outperform baselines in terms of asymptotic performance and sample complexity, enabling tasks beyond current methods? (Q2) How do the key design decisions laid out in Section 3.3 affect performance? (Q3) Can the learned RL policies be used to generate diverse data for sim-to-real transfer?
@@ 4.1 Task Descriptions
4.2 Baseline Comparisons in Simulation
We compare to the following baselines, which bootstrap learning with expert demonstrations, providing them with more prior information about how to solve the task than OmniReset. For each of the methods, we supply 100 successful demonstrations, effectively giving the baselines additional access to the optimal behavior, while OmniReset does not have this information. The initial conditions for these demonstrations are drawn from the reaching region $S^{R}$, which corresponds to the set of initial conditions the robot will encounter when solving the full task.
@@ 4.2 Baseline Comparisons in Simulation
  • 2) DeepMimic: To compare with methods for motion imitation, we use DeepMimic-style reward augmentation (Peng et al., 2018) on top of our generic rewards. During resets, a random demonstration is chosen, and the agent is reset from a random point along the demonstration and is given an auxiliary reward which provides bonuses for following the demonstration.
@@
  • 3) Demo Curriculum: This baseline is constructed in the spirit of method such as (Bauza et al., 2025) and uses a success-weighted autocurriculum to sample reset states from the demonstrations. We use PPO (Schulman et al., 2017) as the base RL algorithm to ensure a fair comparison.
@@
Learning Curves and Success Rates: We plot learning curves for the Hard variants of each task and the Easy variants of the Peg Insertion, Leg Twisting, and Drawer Insertion tasks in Figure 5. Success rates are reported for initial conditions sampled from the Reaching resets $S^{R}$, which correspond to states from which the robot must solve the full task. We see that OmniReset is able to consistently obtain high success rates on each of the tasks, substantially outperforming baseline methods. While the baselines make some progress on the Easy variants of the tasks, they consistently struggle to scale to the wider distributions of initial conditions that define the Hard variants. Appendix Figure 12 provides additional fine-grained insight into the failure modes of the baselines. Here we plot success rates during training for initial conditions drawn from the demonstrations that fall into the Near-Goal and Reaching regions of the state-space. This plot demonstrates how the baselines are able to solve part of the task (i.e, when starting near the goal) but fail to scale to the full long-horizon task (i.e, starting from a reaching position). We refer readers to the supplementary website for a more detailed look at failure modes.
Emergent Curricula: OmniReset does not rely on curricula to stabilize and accelerate learning. We visualize how learning for OmniReset progresses over time on our website. We observe that the diverse resets OmniReset provides enables PPO to naturally solve the problem backwards, first learning to succeed from near-goal states and eventually learning to succeed from the entire search space. This behavior is entirely emergent from our diverse resets and large number of parallel environments.
@@
Robustness of Policies: We conduct a robustness analysis on the learned policies (Fig 6). We sample an initial condition from one of the demonstrations, perturb the initial condition with forces of different magnitudes and report policy success from these perturbed initial conditions. We find that baseline performance quickly degrades under small perturbations, while OmniReset is barely affected under large perturbations.
@@
We also analyze the ability of policies to solve the task from a wide range of initial conditions in the scatter plots in Figure  7. These plots show success rates over various initial conditions for both OmniResetand Demo Curriculum (the most successful baseline) on the Leg Twisting Easy task. For each plot we show the success rate over 1000 sampled initial conditions from the full task distribution. This plot demonstrates how the baseline struggles to achieve achieve consistent success across the distribution of initial conditions it was trained on, while OmniReset achieves is able to succeed across the entire workspace.
Figure 7: Successful initial states of RL policy. For the Lew Twisting, we plot the xy configurations from which RL polices trained with Demo Curriculum and OmniResetsucceed. We find that OmniResetsucceeds from a much broader range of initial conditions.
@@ 4.3 Ablating Key Design Decisions
Figure 8: Ablation on number of environments. We plot the success rates over course of RL using different number of environments. We find that the number of environments significantly impacts training performance.
@@
Figure 9: Ablation on grasp sampling range. For this ablation on the Leg Twisting task, we find that training RL on narrower grasp sampling ranges leads to worse sample efficiency and lower success rate.
@@ 5 Distillation and Real-World Transfer
We demonstrate the utility of our learned data-generation policies by distilling them into visuomotor policies deployable directly on hardware from RGB inputs. Experiments are conducted on a UR7e robot equipped with a Robotiq 2F-85 gripper, with control and policy inference running on a PC with an RTX 4090 GPU. The robot observes $224\times 224$ RGB images from three RealSense cameras: a D455 providing the front view, a D435 for the side view, and a D415 mounted on the wrist. Using the photorealistic rendering capabilities of IsaacLab (Mittal et al., 2023), we collect 80,000 expert rollouts with synchronized images and actions for standard student–teacher distillation Chen et al. (2021). The student policy operates at 10Hz and uses an ImageNet-pretrained ResNet-18 encoder and a Gaussian MLP head conditioned on the five most recent observations. See Appendix A.3 for full implementation details for distillation and transfer.
@@
Visual randomization. As shown in Fig. 11, to mitigate the sim-to-real visual gap, we apply extensive domain randomization following DextrAH-G (Singh et al., 2025), varying lighting, backgrounds, object and robot appearance, and workspace textures. Camera extrinsics are calibrated to the real setup with additional pose and FOV jitter for robustness. We also apply standard image augmentations including color jitter, blur, grayscale, and noise.
Figure 10: Sim2Real Tasks. Simulation setup for pretraining (top) and real-world deployment (bottom) of peg insertion (left), leg twisting (middle), and drawer assembly (right).
@@
Dynamics randomization. To reduce the control gap, we calibrate kinematics to hardware and deploy the same task-space operational space controller (Khatib, 1987) in simulation and reality, with the policy predicting end-effector pose deltas that are converted to torques using identical Jacobian-based control. We perform system identification of key actuator parameters (friction, armature, and delay) to match real joint behavior following PACE (Bjelonic et al., 2025), then randomize controller gains and physical parameters around identified values during RL training. We additionally randomize object mass and friction to improve robustness to contact dynamics, and apply a curriculum that progressively reduces the action space to encourage smaller, smoother motions. We emphasize that this curricula was used to improve sim-to-real transfer, and is not necessary for learning high-performing policies in simulation. Indeed, the simulation only evaluations we report are trained without the curricula, maintaining our claim that OmniReset does not rely on complex training curricula for behavior generation.
Figure 11: Visual randomizations. Examples of domain randomization applied during training. We vary lighting conditions, backgrounds, object and robot appearance, and workspace textures, along with camera pose and field-of-view jitter, to improve robustness to real-world visual variation.
Real-World Transfer. We successfully deploy a distilled OmniReset policy on the Peg Insertion, Leg Twisting, and Drawer Insertion tasks, as shown in Figure 10. The distilled RGB policies, trained on 80,000 simulation trajectories spanning a wide range of initial conditions, achieve zero-shot real-world success rates of 85.37% on Peg, 56.36% on Leg, and 15.38% on Drawer, as summarized in Table 1. This substantially outperforms a behavior cloning Diffusion Policy baseline Chi et al. (2025), which achieves $\sim$2% success across tasks when trained on 100 demonstrations.
We additionally report first-try success, defined as completing the task with a single grasp and execution (i.e., no dropping and regrasping), as well as throughput, measured as successful completions per minute. These metrics highlight that there remains significant room for improvement in efficiency and reliability. Full real-world evaluation details are provided in Appendix A.4.
Qualitatively, the OmniReset visuomotor policy exhibits robust retrying behavior, recovering from initial failures and successfully completing the task. For examples of these behaviors, see our project website. Overall, these results demonstrate that OmniReset provides a scalable foundation for training sim-to-real policies capable of handling substantially broader initial condition distributions than prior approaches.
Table 1: Sim-to-real performance across tasks: We report state-based RL performance in simulation, distilled image policy performance in simulation and real-world deployment, and a behavior cloning baseline trained on only 100 real-world demos.
Metric Peg Leg Drawer
State RL Success Rate (Sim) 94.97% 93.94% 89.60%
Distilled Image Policy Success Rate (Sim) 55.45% 43.69% 84.00%
Distilled Image Policy Success Rate (Real) 85.37% 56.36% 15.38%
Distilled Image Policy First-Try Success Rate (Real) 21.95% 43.64% 5.77%
Distilled Image Policy Throughput (success/min, Real) 1.49 0.63 0.31
Real-Only BC (100 demos) Success Rate (Real) 2.44% 1.82% 1.92%
@@ 6 Conclusion
In this work we presented OmniReset, a simple and scalable system for data generation in simulation for complex, dexterous tasks. The primary insight in OmniReset is showing that a diverse, minimally structured set of reset states paired with large-batch on-policy reinforcement learning in simulation can lead to the emergence of surprisingly complex dexterous behavior. We provide a general purpose recipe to instantiate data generators across a variety of manipulation tasks, and demonstrate both the efficacy of this paradigm in simulation and it’s ability to train robust policies which can be successfully transferred directly to the real world. However, the preset framework has several limitations which leave the door open for exciting future works. OmniReset is dependent on the quality of grasps obtained by the grasp sampler, which can fail to generate diverse grasps on complex, highly non-convex objects. Moreover, pre-computing stable grasps for tasks which require bimanual manipulation or dexterous hands will present additional challenges, and whether the techniques and principles behind OmniReset will scale to such settings remains an open question. Moreover, OmniReset uses relatively modest levels of dynamics randomization for RL training compared existing sim-to-real approaches, and we expect there to be additional challenges training policies which must adapt their behavior to solve the task over a wide range of potential operating conditions. For distillation and transfer, even in simulation our RGB policies achieve much lower success rates than the state-based experts. We expect that reaching higher success rates will require additional research into how to best combine techniques such as DAgger and RL directly from images. Moreover, we found that scaling the RGB dataset to be as large as possible continually improved performance up to the $80k$ trajectories we were able to train on with our compute budget, and we believe obtaining a deeper understanding of scaling laws for this setting will prove valuable. Xu et al. (2023).